OpenStudy (anonymous):

Each week, Rosario drives to an ice skating rink that is 120 miles away. The round-trip takes 2.75 hours. If he averages 55 miles per hour on his way to the rink, which equation can be used to find x, the number of miles per hour he averages on his way home?

OpenStudy (anonymous):

time is distance divided by rate , so time going it \(\frac{60}{55}\) and time returning is \(\frac{60}{x}\) and you know the total is \(2.75\) so \[\frac{60}{55}+\frac{60}{x}=2.75\]

OpenStudy (anonymous):

thank you